# High-order interactions in quantum optomechanics: fluctuations, dynamics and thermodynamics

### v1

@@ IV.1 Three-body interaction
This Hamiltonian can be diagonalized numerically. In particular, in Fig.1a we plot the eigenenergies calculated with respect to both the total Hamiltonian Eq. (15) (blue lines) and the only first-order contribution Eq. (15a) (gray lines) as a function of the mechanical frequency $\Omega$, fixing the coupling at $\epsilon=0.07$. To make this plot, we subtracted the energy of the ground state. It is interesting to observe that the spectrum of Eq. (15a) displays splits of the eigenenergies even when the mechanical frequency is smaller than the optical frequency. Although no terms of the Hamiltonian in Eq. (15a) suggest explicit resonant interactions when $\Omega\leq\omega_{1}$, these energy splits are due to high-order interaction processes between the wall and the cavity mode. Notably, such processes are more effective in the strong coupling regime, where the dynamics become more sensitive to the exchanges of virtual particles [31].
@@
To better understand why such discrepancy between first- and higher-order interactions occurs, we recall that, in case of the first-order resonant interaction, the dominant contribution to the dynamics (and in particular to the boson scattering) is given by those terms of the Hamiltonian which do not rotate in the interaction picture, namely those proportional to $\hat{b}(\hat{a}_{1}^{\dagger})^{2}+\hat{b}^{\dagger}\hat{a}_{1}^{2}$. Since such terms are already included in $\hat{H}_{1}$, namely the interaction Hamiltonian at the lowest order in $\epsilon$, high-order interaction terms such as $\hat{H}_{2}$ and $\hat{H}_{3}$ are not expected to have a decisive impact on the spectrum. On the contrary, the interaction Hamiltonians $\hat{H}_{2}$ and $\hat{H}_{3}$ include high-order resonant terms which play an important role when such high-order resonances are activated.
We now provide a physical interpretation of the mathematical statements made above. The modification of the energy spectrum in proximity to each energy split can be interpreted as the overlap between virtual and real processes that have similar weight in the effective Hamiltonian. Once the first-order resonance has been fixed, the Hamiltonian $\hat{H}_{1}$ explicitly contains Casimir-like terms that permit the conversion of single phonons into real photons. This phenomenon is dominant, since it occurs at the lowest order in $\epsilon$. On the other hand, high-order resonant interactions may take place either via the creation and annihilation of virtual (off-shell) particles predicted in high-order perturbative processes by means of $\hat{H}_{1}$, or due to the direct conversion of phonons into photons enabled by $\hat{H}_{2}$ and $\hat{H}_{3}$.
To further quantify the impact of the nonlinearities on the coupling strength, we can calculate the effective Hamiltonian by means of James’ method [67]. Starting from the Hamiltonian in Eq. (15) and focusing only on the second-order resonance $\Omega=\omega_{1}$, it can be seen that the effective interaction Hamiltonian is $\hat{H}_{\textrm{eff}}=\epsilon^{2}\hat{H}_{\textrm{Kerr}}+\epsilon^{2}\hat{H}_{\textrm{scat}}$, where $\hat{H}_{\textrm{Kerr}}$ includes all Kerr-like frequency shifts, whereas
@@
[…]
IV.2 Dynamics and thermodynamics
The main goal of this section is to examine the effects of high-order interactions encoded in the interaction Hamiltonians $\hat{H}_{2}$ and $\hat{H}_{3}$ on both the dynamics and the thermodynamics of the system. To do this, we study the time-evolution of the system in two different scenarios: in one case, we solve the master equation utilizing the whole Hamiltonian in Eq. (15); in the other, we consider the evolution solely due to Eq. (15a). According to our analysis of the energy spectra, we should expect that the presence of $\hat{H}_{2}$ and $\hat{H}_{3}$ in the Hamiltonian has a higher impact on the dynamics ruled by high-order interactions than on the dynamics first-order one.
IV.2.1 First-order resonance
@@
Figure 5: Time evolution of the populations of the mirror (red line), of the cavity mode 1 (blue line), and of cavity mode 2 (green line), fixing the resonant conditions $\Omega=\omega_{1}$ (a) and $3\Omega=2\omega_{1}$ (b). The solid lines represent the dynamics of the populations evolving via the whole Hamiltonian in Eq. (15), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (15a).
IV.2.2 Second- and third-order resonances
@@
Figure 6: Time evolution of $\mathcal{J}_{\text{w}}(t)$ (red line) and $\mathcal{J}_{\text{c}}(t)$ (blue line) fixing the resonant conditions $\Omega=\omega_{1}$ (a) and $3\Omega=2\omega_{1}$ (b). The solid lines represent the dynamics of the heat flows evolving via the whole Hamiltonian in Eq. (15), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (15a).
More than the phononic population itself, the relevant aspect we want to discuss here is the sensitivity of the phononic population to the high-order Hamiltonian terms. In the previous section we have observed that, once we have fixed the first-order resonance and $\epsilon=0.07$, the presence of $\hat{H}_{2}$ and $\hat{H}_{3}$ (slightly) changes the photonic population, while the phononic population is scarcely affected. Instead, we now observe that the number of phonons changes notably, which seems to be a symptom of a deep modification of the effective coupling between mechanical and optical subsystems. Note that the change of coupling strength is witnessed also by the heat rates and the entropy production rates, plotted in Fig.6 and Fig.7 respectively. Moreover, the dependence of both the number of phonons and photons on the effective coupling, once the resonance condition $\Omega=\omega_{1}$ has been imposed, has already been discussed in the literature [31], where it has also been observed that the increase of the coupling strength induces a higher conversion of hot phonons into photons.
Figure 7: Time evolution of the entropy production rate, fixing the resonant conditions $\Omega=\omega_{1}$ (a) and $3\Omega=2\omega_{1}$ (b). The red solid line represents the dynamics of the heat flows evolving via the whole Hamiltonian in Eq. (15), whereas the black dashed line indicates the evolution due to the only contribution in Eq. (15a).
@@ VI Acknowledgment
Appendix A The interaction Hamiltonians $\hat{H}_{2}$ and $\hat{H}_{3}$
The Hamiltonian calculated in Eq. (III.1) contains parameters that depend on the wave vectors. Since these lengthy expressions do not offer any valuable information, it has been preferred to include them in this appendix. In particular, the interaction sector $\hat{H}_{2}$ contains the function

### v2

@@ III.1 Three-body interaction Hamiltonian
This Hamiltonian can be diagonalized numerically. In particular, in Fig.1a we plot the eigenenergies calculated with respect to both the total Hamiltonian Eq. (8) (blue lines) and the only first-order contribution Eq. (8b) (gray lines) as a function of the mechanical frequency $\Omega$, fixing the coupling at $\epsilon=0.07$. To make this plot, we subtracted the energy of the ground state. It is interesting to observe that the spectrum of Eq. (8b) displays splits of the eigenenergies even when the mechanical frequency is smaller than the optical frequency. Although no terms of the Hamiltonian in Eq. (8b) suggest explicit resonant interactions when $\Omega\leq\omega_{1}$, these energy splits are due to high-order interaction processes between the wall and the cavity mode. Such processes are more effective in the strong coupling regime, where the dynamics becomes sensitive to the exchanges of virtual particles [67].
@@
To better understand the different impact of $\hat{H}_{1}$, $\hat{H}_{2}$ and $\hat{H}_{3}$ on the energy spectrum, we recall that, in case of the first-order resonant interaction $\Omega=2\omega_{1}$, the dominant contribution to the dynamics (and in particular to the boson scattering) is given by those terms of the Hamiltonian that do not rotate in the interaction picture, namely those proportional to $\hat{b}(\hat{a}_{1}^{\dagger})^{2}+\hat{b}^{\dagger}\hat{a}_{1}^{2}$. Such Casimir-like terms permit the conversion of single phonons into real photons, and are already included in $\hat{H}_{1}$, namely the interaction Hamiltonian at the lowest order in $\epsilon$. Any alteration of the spectrum due to the presence of $\hat{H}_{2}$ and $\hat{H}_{3}$ (apart from the Kerr-shift) is negligible. On the contrary, high-order resonant interactions are influenced by all the interaction Hamiltonians $\hat{H}_{1}$, $\hat{H}_{2}$ and $\hat{H}_{3}$. Indeed, fixed $2\Omega=2\omega_{1}$ or $3\Omega=2\omega_{1}$, the photon-phonon conversion takes place via both the creation and annihilation of virtual (off-shell) particles predicted in high-order perturbative processes predicted by $\hat{H}_{1}$, and the conversion of real particles enabled by $\hat{H}_{2}$ and $\hat{H}_{3}$.
The competition of virtual and real processes in high-order resonant interactions has an impact on the effective coupling strength. To quantify it, we can calculate the effective Hamiltonian by means of James’ method [69]. Starting from the Hamiltonian in Eq. (8) and focusing only on the second-order resonance $2\Omega=2\omega_{1}$, it can be seen that the effective interaction Hamiltonian is $\hat{H}_{\textrm{eff}}=\epsilon^{2}\hat{H}_{\textrm{Kerr}}+\epsilon^{2}\hat{H}_{\textrm{scat}}$, where $\hat{H}_{\textrm{Kerr}}$ includes all Kerr-like frequency shifts, whereas
@@
represents the two-phonon-two-photon scattering. The first term of the coupling strength, namely that proportional to $2/\Omega$, stems from virtual processes mathematically described in the interaction picture by the effective Hamiltonian $i\hat{H}_{1}(t)\int\hat{H}_{1}(t^{\prime})dt^{\prime}$ [69]. The second term is already contained in $\hat{H}_{2}$. It is easy to observe that, once the resonance conditions $\omega_{2}=2\omega_{1}=2\Omega$ are imposed, the coupling constant is determined at 71.4% by virtual processes arising from $\hat{H}_{1}$, and at 28.6% by $\hat{H}_{2}$. This implies that more than 1/4 of the coupling strength is given by $\hat{H}_{2}$.
III.2 Dynamics and thermodynamics
In this section, we investigate the heat transfer between the hot bath coupled to the wall and the cold bath coupled to the cavity by analysing the dynamics of the populations of the involved bosonic degrees of freedom, the rates of heat transferred between modes and baths, as well as the entropy production rate. To do so, we study the time-evolution of the system in two different scenarios: in one case, we take into account the whole Hamiltonian in Eq. (8); in the other, we consider the evolution solely due to Eq. (8b). According to our analysis of the energy spectra, we should expect that the presence of $\hat{H}_{2}$ and $\hat{H}_{3}$ in the Hamiltonian has a greater impact on the dynamics ruled by high-order interactions than on the dynamics first-order one.
Results have been obtained by first numerically solving the master equation in Eq. (19) appling numerical filtering technique with parameter $\Delta=0.09$ normalized with respect to $\omega_{2}$ (details on the master equation are discussed in Appendix A.2), and subsequently calculating the average values of each quantity of interest. In particular, thermodynamic quantities such as heat flows and entropy production rate have been defined in Appendix A.3, respectively in Eq. (24) and Eq. (25). For the realization of all graphs we have used the following parameters: $\omega_{2}=2\omega_{1}=1$, $\gamma=0.018$, $\kappa=0.006$, $T_{\textrm{c}}=10^{-6}$ and $T_{\textrm{w}}=0.3$. Moreover, for mere convenience matter, we have chosen $\kappa_{0}=0.003$ as time scale. The quantum mechanical frequency of the wall is determined by the resonant conditions, and it is therefore specified on the caption of each graph. Finally, frequencies and temperatures have been normalized with respect to $\omega_{2}$.
Figure 2: Time evolution of the populations of the mirror (red line), of cavity mode 1 (blue line), and of cavity mode 2 (green line), fixing the resonant condition $\Omega=2\omega_{1}$. The solid lines represent the dynamics of the populations evolving via the whole Hamiltonian in Eq. (8), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (8b). The two graphs are realized by setting $\epsilon=0.03$ (a), and $\epsilon=0.07$ (b).
Figure 3: Time evolution of $\mathcal{J}_{\text{w}}(t)$ (red line) and $\mathcal{J}_{\text{c}}(t)$ (blue line) after enforcing the resonant condition $\Omega=2\omega_{1}$. The solid lines represent the dynamics of the heat flows evolving via the whole Hamiltonian in Eq. (8), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (8b). The graphs are realized by setting $\epsilon=0.03$ (a), and $\epsilon=0.07$ (b).
Figure 4: Time evolution of the entropy production rate fixing the resonant condition $\Omega=2\Omega_{1}$. The solid lines represent the dynamics of the entropy production rate evolving via the whole Hamiltonian in Eq. (8), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (8b). Plots have been realized by setting $\epsilon=0.03$ (red lines), and $\epsilon=0.07$ (blue lines).
III.2.1 First-order resonance
@@
Figure 5: Time evolution of the populations of the mirror (red line), of the cavity mode 1 (blue line), and of cavity mode 2 (green line), fixing the resonant conditions $\Omega=\omega_{1}$ (a) and $3\Omega=2\omega_{1}$ (b). The solid lines represent the dynamics of the populations evolving via the whole Hamiltonian in Eq. (8), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (8b).
III.2.2 Second- and third-order resonances
@@
Figure 6: Time evolution of $\mathcal{J}_{\text{w}}(t)$ (red line) and $\mathcal{J}_{\text{c}}(t)$ (blue line) fixing the resonant conditions $\Omega=\omega_{1}$ (a) and $3\Omega=2\omega_{1}$ (b). The solid lines represent the dynamics of the heat flows evolving via the whole Hamiltonian in Eq. (8), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (8b).
More than the phononic population itself, the relevant aspect we want to discuss here is the sensitivity of the phononic population to the high-order Hamiltonian terms. In the previous section we have observed that, once we have fixed the first-order resonance and $\epsilon=0.07$, the presence of $\hat{H}_{2}$ and $\hat{H}_{3}$ (slightly) changes the photonic population, while the phononic population is scarcely affected. Instead, we now observe that the number of phonons changes notably, which seems to be a symptom of a deep modification of the effective coupling between mechanical and optical subsystems. Note that the change of coupling strength is witnessed also by the heat rates and the entropy production rates, plotted in Fig.6 and Fig.7 respectively. Moreover, the dependence of both the number of phonons and photons on the effective coupling, once the resonance condition $2\Omega=2\omega_{1}$ has been imposed, has already been discussed in the literature [67], where it has also been observed that the increase of the coupling strength induces a higher conversion of hot phonons into photons.
Figure 7: Time evolution of the entropy production rate, fixing the resonant conditions $\Omega=\omega_{1}$ (a) and $3\Omega=2\omega_{1}$ (b). The red solid line represents the dynamics of the entropy production rates evolving via the whole Hamiltonian in Eq. (8), whereas the black dashed line indicates the evolution due to the only contribution in Eq. (8b).
@@ V Acknowledgment
Appendix A Theoretical background
A.1 Calculation of the Hamiltonian
We consider a massless scalar field confined in a 3-D cavity as a simple model for the electromagnetic field to obtain a qualitative understanding of the dynamics. The Lagrangian of the field is
$\mathcal{L}(t,\mathbb{x})=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi.$ (10)
The field confinement is imposed by means of boundary conditions to the equations of motion. If one of the cavity walls oscillates, say along the direction $x$, the wall+field system is described by coupled differential equations for both the wall and the field. However, if the oscillation amplitude of the wall $\delta L_{\text{x}}$ is much smaller than the cavity length $L_{\text{x}}$, i.e., $\delta L_{\text{x}}/L_{\text{x}}\ll 1$, it results much simpler to first assume that the wall is fixed, thereby solving the equation of motion for the field with static Dirichlet boundary conditions, and only subsequently introduce a small fluctuation in the wave vector of the field modes.
To solve the equation of motion, namely the Klein Gordon equation $\partial_{t}^{2}\phi-\nabla^{2}\phi$=0, we consider static Dirichlet boundary conditions. The equation of motion admits as a solution
$\phi(t,\mathbb{x})=$ $\sum_{\mathbb{n}}\left[\alpha_{\mathbb{n}}\,\phi_{\mathbb{n}}(t,x)+\alpha_{\mathbb{n}}^{*}\,\phi^{*}_{\mathbb{n}}(t,x)\right],$ (11)
where each mode $\phi_{\mathbb{n}}(t,\mathbb{x})$ is described by
$\phi_{\mathbb{n}}(t,\mathbb{x})=$ $\sqrt{\frac{4}{\omega_{\mathbb{n}}V}}e^{-i\omega_{\mathbb{n}}t}\chi_{\mathbb{n}}(\mathbb{x})$ (12)
with
$\chi_{\mathbb{n}}(\mathbb{x})=$ $\sin\left(\frac{n_{x}\pi}{L_{x}}x\right)\sin\left(\frac{n_{y}\pi}{L_{y}}y\right)\sin\left(\frac{n_{z}\pi}{L_{z}}z\right)=\sin\left(k_{{n_{\textrm{x}}}}x\right)\sin\left(k_{{n_{\textrm{y}}}}y\right)\sin\left(k_{{n_{\textrm{z}}}}z\right),$ (13)
and it is characterized by the dispersion relation
$\omega_{\mathbb{n}}:=\sqrt{k_{{n_{\textrm{x}}}}^{2}+k_{{n_{\textrm{y}}}}^{2}+k_{{n_{\textrm{z}}}}^{2}},$ (14)
with $\mathbb{n}\equiv(n_{x},n_{y},n_{y})$ a set of positive integer numbers and we have introduced the components of the wave vectors. In Eq. (12), we introduced the volume $V=L_{x}L_{y}L_{z}$.
We can calculate Hamiltonian density $\mathcal{H}(t,\mathbb{x})=\frac{1}{2}\left[\Pi^{2}(t,\mathbb{x})+(\nabla\phi(t,\mathbb{x}))^{2}\right]$, with canonical momentum $\Pi(t,\mathbb{x}):=-\partial_{t}\phi(t,\mathbb{x})$. This reads
$\mathcal{H}(t,\mathbb{x})=$ $\frac{1}{V}\sum_{\mathbb{n}\mathbb{m}}\left[-\sqrt{\omega_{\mathbb{n}}\omega_{\mathbb{m}}}(\alpha_{\mathbb{n}}^{*}-\alpha_{\mathbb{n}})(\alpha_{\mathbb{m}}^{*}-\alpha_{\mathbb{m}})\sin\left(k_{{n_{\textrm{x}}}}x\right)\sin\left(k_{{n_{\textrm{y}}}}y\right)\sin\left(k_{{n_{\textrm{z}}}}z\right)\sin\left(k_{{m_{\textrm{x}}}}x\right)\sin\left(k_{{m_{\textrm{y}}}}y\right)\sin\left(k_{{m_{\textrm{z}}}}z\right)\right$.
$+\frac{k_{{n_{\textrm{x}}}}k_{{m_{\textrm{x}}}}}{\sqrt{\omega_{\mathbb{n}}\omega_{\mathbb{m}}}}(\alpha_{\mathbb{n}}^{*}+\alpha_{\mathbb{n}})(\alpha_{\mathbb{m}}^{*}+\alpha_{\mathbb{m}})\cos\left(k_{{n_{\textrm{x}}}}x\right)\sin\left(k_{{n_{\textrm{y}}}}y\right)\sin\left(k_{{n_{\textrm{z}}}}z\right)\cos\left(k_{{m_{\textrm{x}}}}x\right)\sin\left(k_{{m_{\textrm{y}}}}y\right)\sin\left(k_{{m_{\textrm{z}}}}z\right)$
$+\frac{k_{{n_{\textrm{y}}}}k_{{m_{\textrm{y}}}}}{\sqrt{\omega_{\mathbb{n}}\omega_{\mathbb{m}}}}(\alpha_{\mathbb{n}}^{*}+\alpha_{\mathbb{n}})(\alpha_{\mathbb{m}}^{*}+\alpha_{\mathbb{m}})\sin\left(k_{{n_{\textrm{x}}}}x\right)\cos\left(k_{{n_{\textrm{y}}}}y\right)\sin\left(k_{{n_{\textrm{z}}}}z\right)\sin\left(k_{{m_{\textrm{x}}}}x\right)\cos\left(k_{{m_{\textrm{y}}}}y\right)\sin\left(k_{{m_{\textrm{z}}}}z\right)$
$\left.+\frac{k_{{n_{\textrm{z}}}}k_{{m_{\textrm{z}}}}}{\sqrt{\omega_{\mathbb{n}}\omega_{\mathbb{m}}}}(\alpha_{\mathbb{n}}^{*}+\alpha_{\mathbb{n}})(\alpha_{\mathbb{m}}^{*}+\alpha_{\mathbb{m}})\sin\left(k_{{n_{\textrm{x}}}}x\right)\sin\left(k_{{n_{\textrm{y}}}}y\right)\cos\left(k_{{n_{\textrm{z}}}}z\right)\sin\left(k_{{m_{\textrm{x}}}}x\right)\sin\left(k_{{m_{\textrm{y}}}}y\right)\cos\left(k_{{m_{\textrm{z}}}}z\right)\right]$ (15)
Having obtained the Hamiltonian density, the procedure to quantize the optomechanical system in [21] provides for the following steps:
  1. 1.
    Allow for a small fluctuation of the cavity length by replacing $L_{\text{x}}\rightarrow L_{\text{x}}+\delta L_{\text{x}}$, and then expanding each wave vector $k_{\text{x}}$ in the the Hamiltonian density with respect to the perturbation parameter $\delta L_{\text{x}}/L_{\text{x}}$ up to the third order.
  2. 2.
    Integrate the Hamiltonian density over the box volume.
  3. 3.
    Promote the field mode amplitudes $a_{\mathbb{n}}$ and the small perturbation $\delta L_{\text{x}}$ to the quantum operators $\hat{a}_{\mathbb{n}}$ and $\delta L_{0}(\hat{b}+\hat{b}^{\dagger})$, respectively.
We do not go through the first and the second steps into details, as they are lengthy calculations merely consisting of Maclaurin expansions and integrations over the spatial variables. The last step is complemented by imposing canonical bosonic commutation rules for the oscillation operators $\hat{b},\hat{b}^{\dagger}]=1$ as well as for the field modes $\hat{a}_{\mathbb{n}},\hat{a}_{\mathbb{m}}^{\dagger}]=\delta_{\mathbb{n}\mathbb{m}}$. Note that $\mathbb{n}=(n_{\text{x}},n_{\text{y}},n_{\text{z}})$. We stress that, in this protocol, we expand the Hamiltonian density with respect to $\delta L_{\text{x}}$ only in the wave vectors. This is equivalent to the procedure introduced originally by Law [41], where both the bosonic operators $\hat{a}_{k}$ and the mode frequencies $\omega_{k}$ have been expanded with respect to the oscillation amplitude $x_{m}$ around the average cavity length $l_{0}$.
Once terminated this procedure, we obtain the Hamiltonian in Eq. (1) and Eq. (II.1). In this expression, we defined three functions of the wave vectors, $\zeta_{\mathbb{n}\mathbb{m}}$, $\theta_{\mathbb{n}\mathbb{m}}$ and $\theta_{\mathbb{n}\mathbb{m}}$. In particular, the interaction sector $\hat{H}_{2}$ contains the function
@@ A.1 Calculation of the Hamiltonian
A.2 Quantum state evolution via master equation
Nonlinear interactions introduce an element of anharmonicity, thereby leading to small modifications in the energy level structure while retaining an overall quasi-harmonic character. As a result, the dynamics of the system has to be described using a generalized master equation formulated without invoking the conventional secular approximation. An appropriate theoretical framework for capturing the time evolution of the density matrix operator $\hat{\rho}$ in hybrid quantum systems subject to dissipation and thermal-like noise has been obtained in previous work [68]. The master equation reads $\dot{\hat{\rho}}=-i[\hat{H},\hat{\rho}]+\mathcal{L}\hat{\rho}$, or
$\dot{\hat{\rho}}=-i[\hat{H},\hat{\rho}]+\mathcal{L}_{\textrm{c}}\hat{\rho}+\mathcal{L}_{\textrm{w}}\hat{\rho}\;$, (19)
where we have split the Liouvillian superoperator $\mathcal{L}=\mathcal{L}_{\textrm{c}}+\mathcal{L}_{\textrm{w}}$ in two parts: $\mathcal{L}_{\textrm{c}}$ describes the interaction of the system with the cavity bath at temperature $T_{\textrm{c}}$ and with damping rate $\kappa$, whereas $\mathcal{L}_{\textrm{w}}$ describes the interaction of the system with the wall bath at temperature $T_{\textrm{w}}$ and with damping rate $\gamma$.
The two Liouvillian superoperators are expressed in the form:
$\mathcal{L}_{\textrm{c}}\hat{\rho}=\frac{\kappa}{2}\sum_{\begin{subarray}{c}j>k\\ l>m\end{subarray}}$ $\{A_{lm}n_{ml}({T})[\hat{P}_{lm}\hat{\rho}\hat{P}_{kj}-\hat{P}_{kj}\hat{P}_{lm}\hat{\rho}]+A_{kj}n_{jk}({T})[\hat{P}_{lm}\hat{\rho}\hat{P}_{kj}-\hat{\rho}\hat{P}_{kj}\hat{P}_{lm}]$
$+A_{lm}[n_{ml}({T})+1][\hat{P}_{kj}\hat{\rho}\hat{P}_{lm}-\hat{\rho}\hat{P}_{lm}\hat{P}_{kj}]+A_{kj}[n_{jk}({T})+1][\hat{P}_{kj}\hat{\rho}\hat{P}_{lm}-\hat{P}_{lm}\hat{P}_{kj}\hat{\rho}$]\}\;. (20)
and
$\mathcal{L}_{\textrm{w}}\hat{\rho}=\frac{\gamma}{2}\sum_{\begin{subarray}{c}j>k\\ l>m\end{subarray}}$ $\{B_{lm}n_{ml}({T})[\hat{P}_{lm}\hat{\rho}\hat{P}_{kj}-\hat{P}_{kj}\hat{P}_{lm}\hat{\rho}]+B_{kj}n_{jk}({T})[\hat{P}_{lm}\hat{\rho}\hat{P}_{kj}-\hat{\rho}\hat{P}_{kj}\hat{P}_{lm}]$
$+B_{lm}[n_{ml}({T})+1][\hat{P}_{kj}\hat{\rho}\hat{P}_{lm}-\hat{\rho}\hat{P}_{lm}\hat{P}_{kj}]+B_{kj}[n_{jk}({T})+1][\hat{P}_{kj}\hat{\rho}\hat{P}_{lm}-\hat{P}_{lm}\hat{P}_{kj}\hat{\rho}$]\}\;. (21)
The generalized master equation in Eq. (19) (with the dissipators defined in Eq. (20) and Eq. (21)) is not of Lindblad form, which implies that essential properties such as the positivity of the density matrix and conservation of probability are not automatically ensured. However, a detailed inspection shows that the dissipator can be regarded as approximately lindblad-like under certain conditions [68]. Specifically, if the energy transition rate $|E_{\beta}-E_{\alpha}|\approx|\omega_{\alpha\beta}|$ is much larger than the system damping rates $\gamma,\kappa$, the corresponding terms contribute negligibly to the dynamics. Therefore, one can assume in the Liouvillian superoperator that $n_{jk}({T})\approx n_{ml}({T})$ making the dissipator effectively Lindblad in nature. Despite their limited physical relevance, these quantities result in fast-oscillating terms in the dissipative dynamics that can cause numerical instabilities and significantly increase computational cost.
A.3 Quantum thermodynamics
In the previous section, we presented a method to study the system dynamics using the generalized master equation, under the assumption that $|E_{\beta}-E_{\alpha}|\approx|\omega_{\alpha\beta}|$ much larger than the system damping rates $\gamma,\kappa$. These conditions ensure the positivity of the density matrix and the conservation of probability in numerical simulations. In order to establish an interaction with the environment we assumed that the two subsystems, namely the wall and the cavity field, are coupled to two different baths at different temperatures, $T_{\text{w}}$ and $T_{\text{c}}$ respectively, with $T_{\text{w}}>T_{\text{c}}$ [38]. This suggests that, if the two subsystems also interact with themselves resonantly, heat will flow from the hot bath to the cold bath, and the whole quantum system is constantly out of equilibrium [37, 70]. To describe the dynamics of the heat flows, we first need to recall some fundamental notion of quantum thermodynamics.
In the absence of external drives, i.e. when $\hat{H}$ is time-independent, the first law of thermodynamics in its differential form is recovered by computing the time derivative of the average energy $\langle\hat{H}\rangle:=\textrm{Tr}[\hat{H}\hat{\rho}(t)]$ as follows:
where we have introduced the total heat flow $\mathcal{J}(t)=\mathcal{J}_{\text{c}}(t)+\mathcal{J}_{\text{w}}(t)$. Importantly, we define the heat-flow from the cold bath to the cavity by means of the expression, and the heat-flow from the hot bath to the wall by means of the expression. These are
$\mathcal{J}_{\text{c}}(t)\equiv$ $\dot{\mathcal{Q}}_{\text{c}}(t)=\textrm{Tr}[\hat{H}\hat{\mathcal{L}}_{\text{c}}\hat{\rho}(t)]$ (24a)
$\mathcal{J}_{\text{w}}(t)\equiv$ $\dot{\mathcal{Q}}_{\text{w}}(t)=\textrm{Tr}[\hat{H}\hat{\mathcal{L}}_{\text{w}}\hat{\rho}(t)],$ (24b)
see [37, 36].
$\frac{dS}{dt}$ $=-\mathrm{Tr}\left(\frac{d\hat{\rho}}{dt}\ln\hat{\rho}\right)=-\mathrm{Tr}\left[(\mathcal{L}\hat{\rho})\ln\hat{\rho}\right]=-\mathrm{Tr}[\mathcal{L}_{\mathrm{w}}\hat{\rho}\ln\hat{\rho}]-\mathrm{Tr}[\mathcal{L}_{\mathrm{c}}\hat{\rho}\ln\hat{\rho}$], (26)
$\dot{\Sigma}$ $=-\mathrm{Tr}[\mathcal{L}_{\mathrm{w}}\hat{\rho}\ln\hat{\rho}]-\mathrm{Tr}[\mathcal{L}_{\mathrm{c}}\hat{\rho}\ln\hat{\rho}]-\frac{\mathrm{Tr}(H\mathcal{L}_{\mathrm{w}}\hat{\rho})}{T_{\mathrm{w}}}-\frac{\mathrm{Tr}(H\mathcal{L}_{\mathrm{c}}\hat{\rho})}{T_{\mathrm{c}}}.$ (27)
[…]
where $\alpha=\mathrm{c}$ or $\mathrm{w}$.